a chemist needs 16 liters of a 15% solution. she has two saline solutions with strength 12% and 20%. find the amount of 12% solution and the amount of 20% solution she should add to get 16 liters of a 15% solution.
Answer by richwmiller(17219)

We need a total of 16 liters at 15%
We want to know how many liters at 12% to mix with 20%
a+b=16,
0.12*a+0.2*b=0.15*16
a=16-b
0.12*(16-b)+0.2*b=2.4
1.92-0.12b+0.2*b=2.4
0.08*b=0.48
b=6
a=16-b
a=10 liters at 12%
b=6 liters at 20%
check
0.12*10+0.2*6=0.15*16
1.2+1.2=2.4
2.4=2.4
ok
codemx1